The first part of the shift was dedicated to complete the work on IMC Vertical Accelerometer Recovery and on SIB1 (Boschi, Ruggi).
Then at 17:00 UTC an earthquake of M 6.5 in Afghanistan shaked all the suspensions and opened SIB2 and SNEB; SNEB was easily closed while for SIB2 it was necessary the help of Alessandro.
The ITF recovery started around 18:00; CARM_NULL_1F was achieved at 18:24 UTC. The lock was achived using only two coils of the WI suspension.
Once locked we realized that the structure at 370Hz was very high thus the crew started to investigated on the possible source of this noise; investigation still in progress at the end of the shift.
At 21:56 UTC it was started the ABSORPTION_MEASUREMT.
SUSP
MC horizontal and vertical loop left open.